# Towards Generalizable Autonomous Penetration Testing via Domain Randomization and Meta-Reinforcement Learning

### v1

@@ Abstract
With increasing numbers of vulnerabilities exposed on the internet, autonomous penetration testing (pentesting) has emerged as an emerging research area, while reinforcement learning (RL) is a natural fit for studying autonomous pentesting. Previous research in RL-based autonomous pentesting mainly focused on enhancing agents’ learning efficacy within abstract simulated training environments. They overlooked the applicability and generalization requirements of deploying agents’ policies in real-world environments that differ substantially from their training settings. In contrast, for the first time, we shift focus to the pentesting agents’ ability to generalize across unseen real environments. For this purpose, we propose a Generalizable Autonomous Pentesting framework (namely GAP) for training agents capable of drawing inferences from one to another – a key requirement for the broad application of autonomous pentesting and a hallmark of human intelligence. GAP introduces a Real-to-Sim-to-Real pipeline with two key methods: domain randomization and meta-RL learning. Specifically, we are among the first to apply domain randomization in autonomous pentesting and propose a large language model-powered domain randomization method for synthetic environment generation. We further apply meta-RL to improve the agents’ generalization ability in unseen environments by leveraging the synthetic environments. The combination of these two methods can effectively bridge the generalization gap and improve policy adaptation performance. Experiments are conducted on various vulnerable virtual machines, with results showing that GAP can (a) enable policy learning in unknown real environments, (b) achieve zero-shot policy transfer in similar environments, and (c) realize rapid policy adaptation in dissimilar environments.
Index Terms:
@@
Penetration testing, reinforcement learning, domain randomization, meta-reinforcement learning, large language model.
I Introduction
@@
Penetration testing, shortly pentesting or PT, is an effective methodology to assess cybersecurity through authorized simulated cyber attacks. It aims to preemptively identify security vulnerabilities, allowing organizations to proactively enhance their security measures and defenses. However, with more and more vulnerabilities being exposed to the internet, traditional manual-based pentesting has become more costly, time-consuming, and personnel-constrained [1]. Pentesting is a dynamic sequential decision-making process, where reinforcement learning (RL) emerges as a well-established method for optimizing such decisions. RL trains an agent to learn a policy through trial and error by interacting with the environment, without the need for supervision or predefined environmental models. This makes RL a natural fit for studying autonomous pentesting.
@@
Firstly, training agents in simulated environments is sample-efficient, yet ensuring their realism remains challenging. Emulated environments (e.g., virtual machines) closely mimic real-world settings, while training agents in them is sample-inefficient and time-consuming. To tackle this challenge, we can improve the agents’ learning efficiency in real-world and emulated environments or construct simulated environments close to real-world settings.
@@
Secondly, even with suitable training environments, agents’ learned policies often exhibit poor performance when transferred to unseen testing (real-world) scenarios [3], significantly constraining their widespread applicability. This phenomenon is known as the generalization gap [4, 5]. To tackle this, we need to improve the agents’ generalization ability.
@@
The reasons for the generalization gap are twofold: (a). RL algorithms tend to overfit the training environments [6]; (b). training environments have limited diversity, whereas real-world environments are unknown for agents and unpredictably diverse. The differences between the training and real-world environments are collectively known as the reality gap [7]. As the example depicted in Fig.2, even if vulnerabilities are the same between training and real-world testing environments, the configurations of hosts can differ. For instance, vulnerable products may be exposed on various ports, hosts could be running different operating systems, and there might also be numerous other ports open that are unrelated to the vulnerabilities, as well as potentially different web fingerprints. These differences create a gap between these environments, leading to the agent perceiving different observations. Then, a generalization gap will occur once the agent learns to rely on certain observational features in the training environment that change in the testing environment. We demonstrate this phenomenon in Section V-E, especially in Fig.8.
Figure 1: Gap between the training and testing environments. The training environment is a virtual machine with the CVE-2018-7600 vulnerability, set up using Vulhub [^1] https://github.com/vulhub/vulhub. We utilized the Shodan [^2] https://www.shodan.io engine to search real-world hosts that may possess the same vulnerability, which is assumed to be a testing environment. Even though both training and testing environments have the same vulnerability, differences in host configurations create a gap, thereby affecting the agent’s transfer learning performance.
Previous research on autonomous pentesting mainly focused on improving the agents’ learning performance within specific simulated training environments [8, 9, 10, 11, 12, 13]. They trained agents in highly abstract environments [14, 15, 16], overlooking the challenges related to the applicability and generalization of agents’ policies in real-world settings. In contrast, this paper aims to address the applicability needs of autonomous pentesting, and for the first time, shifts focus to the agents’ ability to generalize across unseen real environments.
@@
We argue that an agent with good generalization ability should be able to draw inferences from one instance to another, akin to human capabilities. Specifically, when the testing environments are similar to the training environment, i.e., have the same vulnerabilities, the agent should be able to bridge the generalization gap and achieve zero-shot policy transfer. Moreover, it should also demonstrate the capacity for few-shot policy adaptation in dissimilar scenarios featuring varying vulnerabilities, thereby enhancing overall learning efficiency.
@@
To this end, we propose a Generalizable Autonomous Pentesting framework, namely GAP. GAP works following a ”Real-to-Sim-to-Real” pipeline. In the Real-to-Sim phase, GAP realizes end-to-end learning in unknown real/emulated environments without requiring significant human effort, while constructing realistic simulation analogs for real environments. On this basis, in the Sim-to-Real phase, GAP aims to enhance agents’ generalization ability, enabling them to achieve two key objectives: (a) bridging the generalization gap for zero-shot policy transfer in similar scenarios, and (b) facilitating rapid policy adaptation in dissimilar scenarios to enhance learning efficiency. For this purpose, two methods are applied:
@@
    We propose GAP, a generalizable autonomous pentesting framework that works on a Real-to-Sim-to-Real pipeline. This framework enables end-to-end policy learning in unknown real environments as well as the construction of realistic simulations, while improving the agent’s generalization ability.
@@
    To bridge the generalization gap and realize fast policy adaption, we are among the first to apply domain randomization in the autonomous pentesting domain and propose an LLM-powered domain randomization method for environment augmentation. We further apply meta-RL to improve the agents’ generalization ability to unseen environments by leveraging the generated simulated training environments.
@@
    We conduct experiments on various vulnerable virtual machines. The experimental results demonstrate that our framework can (a) enable policy learning in unknown real environments, (b) achieve zero-shot policy transfer in similar environments, and (c) realize rapid policy adaptation in dissimilar environments.
II Related Work
II-A Autonomous Penetration Testing
@@
Recently, RL has shown remarkable progress in various domains, including video games [28], autonomous vehicles [29], and notably, cybersecurity [30, 31]. Compared to traditional planning methods, RL allows for training an agent to learn a policy through trial and error while interacting with the environment, eliminating the need for supervision or environmental models. This agent-environment interaction learning paradigm makes RL suitable for studying autonomous pentesting.
Previous research on RL-based autonomous pentesting mainly focused on training agents to discover attack paths in simulated training environments like NASim [16] or CyberBattleSim [14] that are highly abstract and unrealistic. They made attempts to improve the learning efficiency of the pentesting agents through various methods, such as action space decomposition [11], distributed algorithm [32], or the introduction of demonstration data [13] and curiosity mechanism [12]. Remarkable progress has been made, though, they overlooked the generalization and applicability requirements of agents deployed in unknown environments that are different from the training environments.
There also have been some research attempts at applying RL methods in practical pentesting tools [33, 34]. However, these efforts often trained the agent to learn policies in single, fixed scenarios. As a result, such learned policies frequently suffer from poor generalization, as they tend to overfit specific training scenarios or tasks. This limitation hinders their adaptability to the diverse and dynamic network environments encountered in the real world, posing challenges to the widespread deployment of pentesting agents. This motivates us to investigate pentesting agents that can be generalized to unseen real-world scenarios.
II-B Domain Randomization
@@
II-C Meta-Reinforcement Learning
@@
In this paper, we apply MAML to perform meta-training. Compared to recurrence-based methods, gradient-based methods have advantages in parameter efficiency, computational efficiency, and new task adaption. They can also be combined with many policy gradient-based RL algorithms easily since the inner-loop optimization process is just a policy gradient algorithm [45]. Furthermore, under certain assumptions, gradient-based methods can enhance policy performance from the initialization on any given task, even one from outside the task distribution [22].
III Preliminary
III-A Reinforcement Learning
@@
III-B Meta-Reinforcement Learning
@@
III-C Generalization in Reinforcement Learning
@@
We evaluate the agents’ generalization ability in two ways. Firstly, following [5, 48], we first study the zero-shot policy transfer performance in testing environments using the generalization gap, which is defined as
@@
where $G(\tau)$ is the expected cumulative reward over a trajectory, and $\mathbb{E}_{\mathcal{M}_{test},\tau\sim P_{\pi}}\left[G(\tau)\right]$ is the zero-shot performance. This metric measures the agent’s ability to overcome overfitting and achieve zero-shot generalization. A smaller generalization gap indicates that the deployment performance won’t deviate as much from the training performance, potentially indicating a more robust policy. Besides, following [5, 49], we also study the few-shot policy adaption performance by allowing the agent to interact with the testing environment online. This metric can evaluate how well the agent can quickly adapt its learned policy to new tasks or environments with stronger forms of variation.
IV Proposed Method
IV-A Overview
GAP aims to train the pentesting agent to realize drawing inferences from one instance to another, where ”learning from one instance” is the foundation skill, and ”generalizing to another” is the core objective. To achieve this goal, it employs a Real-to-Sim-to-Real pipeline that integrates end-to-end policy learning, the construction of realistic simulation environments, and the enhancement of the agent’s generalization ability. The full workflow of GAP is illustrated in Fig. 2.
@@ IV-A Overview
IV-A1 Real-to-Sim
The increasing emergence of vulnerabilities, coupled with the complex and dynamic nature of real-world network environments, necessitates the ability of pentesting agents to learn and adapt policies in an end-to-end manner within unknown settings. GAP achieves this through the Real-to-Sim stage (see Section IV-B for details), as shown by the blue lines in Fig.2, that is: ① end-to-end policy learning in the original training environment, and ② realistic simulation construction based on information gathered during policy learning. This constructed simulation serves as a digital mapping of the real environment, facilitating efficient policy training and environment augmentation in the subsequent stage.
IV-A2 Sim-to-Real
Sim-to-Real is a comprehensive concept that has been applied to robotic and classic machine vision tasks [50]. The goal of sim-to-real transfer is to ensure that the behaviors, actions, or decisions learned in simulation can effectively and reliably be applied in real-world scenarios. Similarly, in the Sim-to-Real stage, GAP utilizes the simulated environment constructed in the last stage to improve the agent’s generalization ability, allowing the agent to quickly adapt to diverse real-world settings. As shown by the red lines in Fig.2, GAP achieves this through two-phase steps: ③ synthetic environment generation via LLM-powered domain randomization (see Section IV-C), as well as ④ meta-RL training based on the synthetic environments (see Section IV-D). In this way, the agent is expected to achieve zero-shot policy transfer in similar environments (step ⑤ in Fig.2) and few-shot policy adaption in dissimilar environments (step ⑥ in Fig.2).
@@ IV-A2 Sim-to-Real
IV-B Policy Learning and Simulation Construction
@@
IV-B1 Model penetration testing as an MDP
To train the pentesting agent using RL algorithms, we should model the pentesting process as an MDP. This is formalized by defining the state space $\mathcal{S}$, action space $\mathcal{A}$, and reward function $\mathcal{R}$. For model-free RL algorithms, the transition probability function $\mathcal{P}$ remains unknown.

### v2

# Mind the Gap: Towards Generalizable Autonomous Penetration Testing via Domain Randomization and Meta-Reinforcement Learning

@@ Abstract
With increasing numbers of vulnerabilities exposed on the internet, autonomous penetration testing (pentesting) has emerged as a promising research area. Reinforcement learning (RL) is a natural fit for studying this topic. However, two key challenges limit the applicability of RL-based autonomous pentesting in real-world scenarios: (a) training environment dilemma – training agents in simulated environments is sample-efficient while ensuring their realism remains challenging; (b) poor generalization ability – agents’ policies often perform poorly when transferred to unseen scenarios, with even slight changes potentially causing significant generalization gap. To this end, we propose GAP, a generalizable autonomous pentesting framework that aims to realizes efficient policy training in realistic environments and train generalizable agents capable of drawing inferences about other cases from one instance. GAP introduces a Real-to-Sim-to-Real pipeline that (a) enables end-to-end policy learning in unknown real environments while constructing realistic simulations; (b) improves agents’ generalization ability by leveraging domain randomization and meta-RL learning. Specially, we are among the first to apply domain randomization in autonomous pentesting and propose a large language model-powered domain randomization method for synthetic environment generation. We further apply meta-RL to improve agents’ generalization ability in unseen environments by leveraging synthetic environments. The combination of two methods effectively bridges the generalization gap and improves agents’ policy adaptation performance. Experiments are conducted on various vulnerable virtual machines, with results showing that GAP can enable policy learning in various realistic environments, achieve zero-shot policy transfer in similar environments, and realize rapid policy adaptation in dissimilar environments. The source code is available at https://github.com/Joe-zsc/GAP.
keywords:
@@
Cybersecurity, Penetration testing, Reinforcement learning, Domain randomization, Meta-reinforcement learning, Large language model.
journal: xxx
[…]
[inst1]organization=College of Electronic Engineering, National University of Defense Technology,city=Hefei, postcode=230037, country=China
[…]
1 Introduction
@@
Penetration testing, shortly pentesting or PT, is an effective methodology to assess cybersecurity through authorized simulated cyber attacks. It aims to preemptively identify security vulnerabilities, allowing organizations to proactively enhance their security measures and defenses. However, with more and more vulnerabilities exposed on the internet, traditional manual-based pentesting has become more costly, time-consuming, and personnel-constrained [1]. In light of this, autonomous pentesting has emerged as a promising research area. Pentesting is a dynamic sequential decision-making process, while reinforcement learning (RL) is a suitable method for optimizing such decisions. RL trains an agent to learn a policy through trial and error by interacting with the environment, without the need for supervision or predefined environmental models. This makes RL a natural fit for studying autonomous pentesting.
@@
Challenge 1: training environment dilemma. Training agents in simulated environments is sample-efficient, yet ensuring their realism remains challenging. For instance, previous research in RL-based autonomous pentesting [3, 4, 5, 6, 7, 8, 9, 10] mainly focused on training agents to find attack paths within simulated environments like NASim [11] or CyberBattleSim [12], while these environments are highly abstract and unrealistic as they assume that the network environment can be observed from a global perspective before training, which is unrealistic for pentesting in the real world [13]. By contrast, emulated environments (e.g., virtual machines) closely mimic real-world settings, but training agents in them is sample-inefficient and time-consuming. To tackle this challenge, one can improve the agents’ learning efficiency in emulated environments or construct simulated environments close to real-world settings.
@@
Challenge 2: poor generalization ability. That is, even with suitable training environments, agents’ learned policies often exhibit poor performance when transferred to unseen testing (real-world) scenarios [14]. The worst of it is even a slight change in the environment would require the agent’s policy to be retrained. This phenomenon is known as the generalization gap [15, 16], which significantly constrains the widespread applicability of pentesting agents. To tackle this, one needs to improve agents’ generalization ability. The reasons for the generalization gap are twofold:
  1. 1.
    RL algorithms tend to overfit the training environments [17];
  2. 2.
    Training environments have limited diversity, whereas real-world environments are unknown for agents and unpredictably diverse. The differences between the training and real-world environments are collectively known as the reality gap [18].
@@
Fig.2 presents an example of the reality gap and generalization gap encountered in the pentesting field. As the example depicted in Fig.1a, even if vulnerabilities are the same between training and real-world testing environments, the configurations of hosts can differ. For instance, vulnerable products may be exposed on various ports, hosts could be running different operating systems, and there might also be numerous other ports open that are unrelated to the vulnerabilities, as well as potentially different web fingerprints. These differences create a gap between these environments, leading to the agent perceiving different observations. Then, a generalization gap will occur once the agent learns to rely on certain observational features in the training environment that change in the testing environment. Fig. 1b visually demonstrates the impact of the generalization gap on the agent’s policy transfer. We train the agent using April [13] (the state-of-the-art framework for autonomous pentesting) in the training environment shown in Fig. 1a, and then transfer the trained agent to the testing environment for evaluation. The zero-shot performance in the testing environment shows a significant decline compared to the learning performance in the training environment. We further demonstrate this phenomenon in Section 6.2, especially in Fig.8.
@@
We argue that an agent with good generalization ability should be able to draw inferences about other cases from one instance, akin to human capabilities. Specifically, when the testing environments are similar to the training environment, i.e., have the same vulnerabilities, the agent should be able to bridge the generalization gap and achieve zero-shot policy transfer. Moreover, it should also demonstrate the capacity for few-shot policy adaptation in dissimilar scenarios featuring varying vulnerabilities, thereby enhancing overall learning efficiency.
@@
To tackle the two challenges, we propose a Generalizable Autonomous Pentesting framework, namely GAP. GAP works following a ”Real-to-Sim-to-Real” pipeline, where the training environment dilemma and poor generalization ability can be tackled. In the Real-to-Sim phase, GAP realizes end-to-end learning in unknown real/emulated environments without requiring significant human effort, while constructing realistic simulation analogs for real environments. On this basis, in the Sim-to-Real phase, GAP aims to enhance agents’ generalization ability, enabling them to achieve two key objectives: (a) bridging the generalization gap for zero-shot policy transfer in similar scenarios, and (b) facilitating rapid policy adaptation in dissimilar scenarios to enhance learning efficiency. For this purpose, two methods are applied:
@@
    We propose GAP, a generalizable autonomous pentesting framework that works on a Real-to-Sim-to-Real pipeline. This framework enables end-to-end policy learning in unknown real environments as well as the construction of realistic simulations, while improving agents’ generalization ability.
@@
    To bridge the generalization gap and realize fast policy adaptation, we are among the first to apply domain randomization in the autonomous pentesting domain and propose an LLM-powered domain randomization method for environment augmentation. We further apply meta-RL to improve the agents’ generalization ability to unseen environments by leveraging the generated simulations.
@@
    We conduct experiments on various vulnerable virtual machines. The experimental results demonstrate that our framework can (a) enable policy learning in various realistic environments, (b) achieve zero-shot policy transfer in similar environments, and (c) realize rapid policy adaptation in dissimilar environments.
(a) Reality gap: hosts with the same vulnerability have different configurations.
(b) Generalization gap between training and testing environments.
Figure 1: Gap between the training and testing environments. (a) Shows the reality gap between training and testing environments. The training environment is a virtual machine with the CVE-2018-7600 vulnerability, set up using Vulhub [^1] https://github.com/vulhub/vulhub. We utilized the Shodan [^2] https://www.shodan.io engine to search real-world hosts that may possess the same vulnerability, which is assumed to be a testing environment. Even though both training and testing environments have the same vulnerability, differences in host configurations create a gap, thereby affecting the agent’s transfer learning performance. (b) Shows the learning curve of April in the training environment and its zero-shot performance in the testing environment. The averages and 95% confidence intervals (shaded areas) across 3 random seeds are shown.
2 Related Work
2.1 Autonomous Penetration Testing
@@
Recently, RL has shown remarkable progress in various domains, including video games [30], autonomous vehicles [31], unmanned aerial vehicles [32], and cybersecurity [33]. Compared to traditional planning methods, RL allows for training an agent to learn a policy through trial and error while interacting with the environment, eliminating the need for supervision or environmental models. This agent-environment interaction learning paradigm makes RL suitable for studying autonomous pentesting.
Previous research on RL-based autonomous pentesting mainly focused on training agents to discover attack paths in simulated training environments like NASim [11] or CyberBattleSim [12] that are highly abstract and unrealistic. They made attempts to improve the learning efficiency of the pentesting agents through various methods, such as action space decomposition [6], distributed algorithm [34], or the introduction of demonstration data [8] and curiosity mechanism [35]. Besides, [9] and [10] have provided solutions to enhance the adaptability of agents in response to dynamic changes in such simulated environments. Remarkable progress has been made, though, they heavily rely on an over-optimistic and unrealistic assumption that the network environment, including topologies, host configurations, and even vulnerabilities, can be observed from a global perspective, which is unrealistic even for human experts [13].
There has also been some research [36, 37, 13] attempts to train pentesting agents in more realistic environments. However, these efforts often trained agents to learn policies in fixed scenarios. Such learned policies frequently suffer from poor generalization, as they tend to overfit specific training scenarios or tasks. Consequently, even a slight change in the environment would require agents’ policies to be retrained. This limitation hinders their adaptability to the diverse vulnerability environments encountered in the real world, posing challenges to the widespread deployment of pentesting agents.
Different from previous research, this work follows the settings of [13], with the goal of training pentesting agents that can be generalized to unseen real-world scenarios. This generalization ability refers to drawing inferences from one instance to another, enabling agents to bridge the generalization gap for zero-shot policy transfer in similar environments with the same vulnerabilities, as well as achieve few-shot policy adaptation in dissimilar scenarios.
2.2 Domain Randomization
@@
2.3 Meta-Reinforcement Learning
@@
In this paper, we apply MAML to perform meta-training. Compared to recurrence-based methods, gradient-based methods have advantages in parameter efficiency, computational efficiency, and new task adaptation. They can also be combined with many policy gradient-based RL algorithms easily since the inner-loop optimization process is just a policy gradient algorithm [48]. Furthermore, under certain assumptions, gradient-based methods can enhance policy performance from the initialization on any given task, even one from outside the task distribution [24].
3 Preliminary
3.1 Reinforcement Learning
@@
3.2 Meta-Reinforcement Learning
@@
3.3 Generalization in Reinforcement Learning
@@
We evaluate the agents’ generalization ability in two ways. Firstly, following [16, 51], we study the zero-shot policy transfer performance in testing environments using the generalization gap, which is defined as
@@
where $G(\tau)$ is the expected cumulative reward over a trajectory, $\mathbb{E}_{\mathcal{M}_{test},\tau\sim P_{\pi}}\left[G(\tau)\right]$ is the zero-shot performance. This metric measures the agent’s ability to overcome overfitting and achieve zero-shot generalization. A smaller generalization gap indicates that the deployment performance won’t deviate as much from the training performance, potentially indicating a more robust policy. Besides, following [16, 52], we also study the few-shot policy adaptation performance by allowing the agent to interact with the testing environment online. This metric can evaluate how well the agent can quickly adapt its learned policy to new tasks or environments with stronger forms of variation.
4 Proposed Method
4.1 Overview
GAP aims to realize efficient policy training in realistic environments and train generalizable agents capable of drawing inferences about other cases from one instance. ”Learning from one instance” is the foundational skill, while ”generalizing to another” is the core objective. To achieve this goal, it employs a Real-to-Sim-to-Real pipeline that integrates end-to-end policy learning, the construction of realistic simulation environments, and the enhancement of the agent’s generalization ability. The full workflow of GAP is illustrated in Fig. 2.
@@ 4.1 Overview
4.1.1 Real-to-Sim
The increasing emergence of vulnerabilities, coupled with the complex and dynamic nature of real-world network environments, necessitates the ability of pentesting agents to learn and adapt policies in an end-to-end manner within unknown settings. GAP achieves this through the Real-to-Sim stage (see Section 4.2 for details), as shown by the blue lines in Fig.2, that is: ① end-to-end policy learning in the original training environment, and ② realistic simulation construction based on information gathered during policy learning. This constructed simulation serves as a digital mapping of the real environment, facilitating efficient policy training and environment augmentation in the subsequent stage.
4.1.2 Sim-to-Real
Sim-to-Real is a comprehensive concept that has been applied to robotic and classic machine vision tasks [53]. The goal of sim-to-real transfer is to ensure that the behaviors, actions, or decisions learned in simulation can effectively and reliably be applied in real-world scenarios. Similarly, in the Sim-to-Real stage, GAP utilizes the simulated environment constructed in the last stage to improve the agent’s generalization ability, allowing the agent to quickly adapt to diverse real-world settings. As shown by the red lines in Fig.2, GAP achieves this through two-phase steps: ③ synthetic environment generation via LLM-powered domain randomization (see Section 4.3), as well as ④ meta-RL training based on the synthetic environments (see Section 4.4). In this way, the agent is expected to achieve zero-shot policy transfer in similar environments (step ⑤ in Fig.2) and few-shot policy adaptation in dissimilar environments (step ⑥ in Fig.2).
@@ 4.1.2 Sim-to-Real
4.2 Policy Learning and Simulation Construction
@@
4.2.1 Model penetration testing as an MDP
To train the pentesting agent using RL algorithms, we model the pentesting process as an MDP following [13]. This is formalized by defining the state space $\mathcal{S}$, action space $\mathcal{A}$, and reward function $\mathcal{R}$. For model-free RL algorithms, the transition probability function $\mathcal{P}$ remains unknown.